OpenStudy (anonymous):

Can someone check my answer please? Do the values in the table represent a linear function? if so, what is the function rule? (A)the values do not show a linear function (B)Yes, the values show a linear function: y=2x+2 (C)Yes, the values show a linear function: y=1/2x+4. (D)yes, the values show a linear function: y=2x. I pick C but I'm not sure?

OpenStudy (anonymous):

Here is the graph

OpenStudy (loser66):

Nope, not that x =-2, y =-4 x =0, y =0 x =2 , y = 4 x =4, y =8 that shows y =2 x , not 1/2 x =4 if y = (1/2)x +4, then when x =0 , y =4, not =0 as the table showed

OpenStudy (anonymous):

Yeah thank you I already fixed it.